# Sliding DFT-based Signal Recovery for Modulo ADC with 1-bit Folding Information

### v1

@@ Abstract
The modulo analog-to-digital converter (ADC) is a promising solution to resolve the limited dynamic range (DR) issue of conventional ADCs. However, a modulo ADC requires an unfolding scheme to correct the nonlinear distortion introduced by the modulo operation. This paper presents a sliding discrete Fourier Transform (DFT)-based method for fast signal reconstruction given the modulo ADC output sequence and a 1-bit folding information sequence. In contrast to existing DFT-based signal recovery techniques for modulo ADCs, our proposed sliding DFT method reduces the required observation time and minimizes the spectral leakage effects via proper choice of window function parameters. A mean squared error (MSE) performance guarantee is established for the proposed signal recovery algorithm. More precisely, we derive sufficient conditions for the oversampling factor ($\mathrm{OF}$) and the number of quantization bits ($b$) to obtain a specific MSE performance. Our numerical results demonstrate that modulo ADCs equipped with our proposed recovery method can outperform conventional ADCs without modulo for $\mathrm{OF}\geq 4$ and $b\geq 4$. The impact of spectral leakage on the MSE performance of the proposed sliding DFT recovery method is also quantified.
@@
Modulo ADC, Sampling, Quantization, DFT.
@@ I Introduction
A promising solution to address the DR bottleneck of ADCs is the modulo sampling framework [6]. The idea in modulo sampling is to apply a nonlinear folding operation called modulo to the input signal before the ADC. This modulo preprocessing stage ensures that the signal range is confined within the DR of the ADC. Since modulo operation intentionally introduces nonlinear distortion to the signal, several studies on modulo sampling have focused on the development of robust recovery techniques to correct the distortion induced by the modulo operation. Reconstruction techniques for modulo sampling include higher-order differences-based approaches [6], discrete Fourier transform (DFT)-based techniques [7, 8, 9, 10, 11], prediction-based filtering [12], iterative methods [13], sparse signal modulo recovery [14, 15, 16], and dual-channel modulo structures [17]. There are also research efforts geared toward the practical hardware implementation of modulo ADCs [18, 19, 20] and their potential applications in radars, communication systems, imaging, and electroencephalogram (EEG) recovery [21, 22, 23, 24, 25, 26, 27, 17, 28].
@@
Another important research direction in modulo ADC is the establishment of theoretical performance guarantees to reconstruct the unfolded signal from modulo samples. In [6], it is shown that the higher-order differences-based reconstruction approach can perfectly reconstruct a finite-energy bandlimited (BL) signal from its modulo samples if the sampling rate used is at least $2\pi e$ ($\approx 17.07$) times the Nyquist rate. This sufficient condition for the sampling rate increases significantly in the presence of bounded noise [6, Theorem 3]. For finite-energy BL signals, perfect reconstruction is possible for any sampling rate approaching the Nyquist rate from above by using a prediction filter with sufficiently long filter length [20]. However, the proposed approach did not take into account the impact of finite quantization. Theoretical guarantees for Fourier-based reconstruction had been established in [7]. However, this approach suffers from spectral leakage and may require long observation windows. A recent work on modulo ADCs with 1-bit side information [10] established a mean squared error (MSE) performance guarantee when the oversampling factor (OF) is above three and the number of quantization bits is above three. The 1-bit side information identifies the indices in which folding occured. However, the proofs neglected the spectral leakage effects and the contribution of quantization noise at certain time indices.
@@
In this work, we propose a sliding DFT approach to construct the original input samples from the modulo ADC output samples and an additional 1-bit folding information signal. The recovery procedure uses short-time Fourier transform (STFT) to facilitate successive frame-based unfolding of the modulo samples. To significantly reduce the impact of spectral leakage in STFT computation, windowing is incorporated in the recovery method. A MSE performance guarantee for the proposed recovery method is also established and is compared with the simulated MSE results.
@@
  • We develop an unfolding algorithm for modulo ADC output with 1-bit folding information signal based on sliding-window DFT. In contrast to existing Fourier-based reconstruction techniques, the proposed algorithm does not need a long observation window to process the modulo ADC output. This significantly reduces the computational complexity of the recovery algorithm. Furthermore, spectral leakage due to the short frame observations can be reduced by choosing an appropriate window function.
  • We establish sufficient conditions for the oversampling factor (OF) and quantizer resolution of the modulo ADC so that the proposed recovery algorithm can correctly unfold the modulo ADC output. Under these conditions, the MSE performance of the sliding DFT recovery algorithm is derived (see Theorem 1).
@@
Notation: The following notations are used throughout the paper. The sets of real numbers, integers, and natural numbers are denoted as $\mathbb{R}$, $\mathbb{Z}$, and $\mathbb{N}$, respectively. When referring to discrete-time signals, the notation $z[n]$ is used to signify $z(nT_{\mathrm{s}})$, assuming that the sampling period $T_{\mathrm{s}}$ is evident from the context. The first-order difference of a discrete-time signal $x[n]$ is denoted as $\underline{x}[n]=x[n]-x[n-1]$, where $x[-1]=0$ unless stated otherwise. Vectors and matrices are written in bold format (e.g., $\mathbf{z}$, $\mathbf{A}$) while sets are written in calligraphic format (e.g., $\mathcal{S}$). If $\mathbf{x}$ is a vectorized form of some discrete-time signal $x[n]$, then $\underline{\mathbf{x}}$ is the vectorized form of $\underline{x}[n]$. The $\ell_{p}$-norm of a vector $\mathbf{x}$ is written as $\|\mathbf{x}\|_{p}$. The set of values of the discrete-time signal $x[n]$ for $n\in\mathcal{S}$ is denoted as $x_{\mathcal{S}}[n]$. The Moore-Penrose inverse of a matrix $\mathbf{A}\in\mathbb{R}^{m\times n}$ is denoted by $\mathbf{A}^{\dagger}$. We also use the standard Big-O notation $\mathcal{O}(\cdot)$ to describe asymptotic growth rates in this paper.
@@ II-A Signal Acquisition Process
Figure 1: The schematic diagram of the modulo ADC with 1-bit side information. The 1-bit level crossing information $c[n]$ is obtained by sampling $c(t)$.
Figure 2: The folding mechanism and side information generation of the modulo operator block.
We consider the modulo ADC system shown in Figure 1. The input $f(t)$ is a bandlimited signal with (angular) frequency support $\left[-\frac{\omega_{m}}{2},\;+\frac{\omega_{m}}{2}\right]$. This signal is folded by the modulo operator block to produce the modulo signal $f_{\lambda^{\prime}}(t)$. The operation behind the modulo folding mechanism is shown in Figure 2. This feedback structure for the folding operation is adopted from [18]. Comparator 1 triggers a positive value whenever $f_{\lambda^{\prime}}(t)$ crosses $+\lambda$ from below. Similarly, comparator 2 triggers a negative value whenever $f_{\lambda^{\prime}}(t)$ crosses $-\lambda$ from above. These trigger signals are fed to a discrete voltage generator (DVG). The voltage level of the DVG is increased (resp. decreased) by $2\lambda^{\prime}$ whenever it receives a positive (resp. negative) trigger signal. A detailed circuit-level implementation of the DVG is presented in [18]. Mathematically, the folded signal can be expressed as
@@
where $\lambda^{\prime}\in(0,\|f(t)\|_{\infty})$ is the modulo threshold. The $\ell_{\infty}$-norm of a waveform is its maximum amplitude.
To digitize the modulo signal, $f_{\lambda^{\prime}}(t)$ is first sampled at every $T_{s}$ seconds. The angular sampling rate is $\omega_{s}=\frac{2\pi}{T_{s}}=\mathrm{OF}\times\omega_{m}$, where $\mathrm{OF}\geq 1$ is the oversampling factor. We also define $\rho=\frac{1}{\mathrm{OF}}$. Prior to quantization, a (non-subtractive) dither sequence $d[n]$ is added to $f_{\lambda^{\prime}}[n]$. Samples of the dither sequence are drawn i.i.d. from a triangle distribution with amplitude support $\big{(}-\frac{2\lambda}{2^{b}},+\frac{2\lambda}{2^{b}}\big{]}$. The rationale for using triangle dither in the modulo ADC problem setup is explained in Section III-B. The resulting signal is then forwarded to the $b$-bit uniform scalar quantizer $\mathcal{Q}_{b}(\cdot)$ with range $[-\lambda,+\lambda]$. Consequently, each quantization bin has width $\frac{2\lambda}{2^{b}}$. To prevent the quantizer from being overloaded, the modulo threshold is set to $\lambda^{\prime}=\frac{(2^{b}-2)\lambda}{2^{b}}$. The quantizer output can be written as
@@
In addition to $f_{\lambda^{\prime}}(t)$, the modulo operator also generates a 2-level signal $c(t)$ according to Figure 2. This signal is a finite rate of innovation (FRI) signal and can be expressed as
@@
where $\tau_{k}$ is the location of the level crossings, $g_{1}(t)$ is the pulse shape of the trigger signals, $g_{2}(t)$ is the sampling kernel, and $g(t)=(g_{2}\ast g_{1})(t)$. To simplify the analysis, $g(t)$ is assumed to be a rectangular pulse with width $T_{s}$ and unity amplitude. The signal $c(t)$ is sampled at rate $\omega_{s}$ and then fed to a 1-bit comparator to produce $c[n]$. In essence, $c[n]$ provides information if the signal crossed either $+\lambda^{\prime}$ or $-\lambda^{\prime}$ in the interval $(nT_{\mathrm{s}},(n+1)T_{\mathrm{s}}]$.
@@ II-B Is the 1-bit Folding Information Practical?
One might wonder if generating 1-bit folding information $c[n]$ is practical when implementing a modulo ADC. Early works on modulo ADC hardware have considered recovery methods that utilize the reset count in addition to the modulo samples [29, 30]. However, implementing a modulo ADC with a multi-bit reset count signal may incur a significant increase in power consumption and circuit complexity if the modulo threshold is small [6]. Limiting $c[n]$ to 1-bit solves these problems. As shown in Figures 1 and 2, generating $c[n]$ would only require one additional comparator and one adder from the “no $c[n]$” modulo ADC setup. The additional circuitry has minimal impact on the overall power consumption, and its penalty is fixed regardless of $b$. For instance, an 8-bit modulo ADC with no $c[n]$ would have $2582^{8}+2=$ 258 comparators [^1] The 2 additional comparators come from the comparators of the folding mechanism. whereas an 8-bit modulo ADC with 1-bit folding information $c[n]$ would have 259 comparators. We also note that the folding information generated in Figure 1 is one bit less than the folding information generated by the modulo ADC hardware proposed in [19] [^2] In [19], the 2-bit signal indicates three states: ‘no crossing’, ‘crossed $+\lambda^{\prime}$ threshold’, ‘crossed $-\lambda^{\prime}$ threshold’..
@@ II-C Signal Recovery Scheme
$=$ $w[n]\cdot\left(f_{\lambda^{\prime},\mathrm{q}}^{(i)}[n]-f_{\lambda^{\prime},\mathrm{q}}^{(i)}[n-1]\right).$ (3)
We now describe the computation of the modulo residue samples in $f_{\lambda^{\prime},\mathrm{q}}^{(i)}[n]$. The length-$N$ Discrete Fourier Transform (DFT) of $\underline{\tilde{f}_{w}^{(i)}}[n]$ is computed as
@@
$\hat{F}_{\mathrm{OOB}}(e^{j\frac{2\pi k}{N}})=$ $\frac{1}{\sqrt{N}}\sum_{n=0}^{L-1}w[n]\bigg{[}\underline{z^{(i)}}[n]+\underline{\epsilon^{(i)}}[n]\bigg{]}e^{-j\frac{2\pi kn}{N}}$
@@
where the spectral leakage width, denoted $\Omega_{\mathrm{SL}}$, accounts for the frequencies where spectral leakage is still significant. Some energy of the $\underline{\tilde{f}_{w}^{(i)}}[n]$ may leak in the band $(\rho\pi,2\pi-\rho\pi)$ due to the inherent spectral leakage in the DFT computation and cannot be neglected. However, spectral leakage can be controlled by properly choosing the window length $N$ and roll-off parameter $\alpha$. The signals $\underline{z^{(i)}}[n]$ and $\underline{\epsilon^{(i)}}[n]$ are the first-order difference of the modulo residue samples and quantization noise, respectively, in the $\tilde{f}_{w}^{(i)}[n]$. The set $\mathcal{S}_{i}$ contains the indices for which $c[n]=1$ in the $i$-th segment. The set identifies the indices in the $i$-th segment for which the DVG changes its output voltage. The samples $z^{(i)}_{\mathcal{S}_{i}}[n]$ (resp. $z^{(i)}_{\mathcal{S}_{i}^{c}}[n]$) corresponds to the value of the modulo residue for all $n$ such that $c[n]=1$ (resp. $c[n]=0$) in the $i$-th segment. The subscript $\mathrm{OOB}$ in $\hat{F}_{\mathrm{OOB}}$ is used to indicate that this is the set of DFT coefficients outside the desired signal bandwidth plus non-negligible spectral leakage.
We are primarily interested in estimating $z^{(i)}_{\mathcal{S}_{i}}[n]$. Let $K$ be the number of discrete frequencies $k$ such that $\frac{2\pi k}{N}\in(\rho\pi+\Omega_{\mathrm{SL}},2\pi-\rho\pi-\Omega_{\mathrm{SL}})$. Equation (II-C) can be written in matrix form as
@@
where $\underline{\mathbf{z}_{w}}\in\mathbb{R}^{N\times 1}$, $\underline{\boldsymbol{\epsilon}_{w}}\in\mathbb{R}^{N\times 1}$, and $\hat{\mathbf{F}}_{\mathrm{OOB}}\in\mathbb{C}^{K\times 1}$ are vectorized form of the finite-length signals $w[n]\cdot\underline{z^{(i)}}[n]$, $w[n]\cdot\underline{\epsilon^{(i)}}[n]$, and $\hat{F}_{\mathrm{OOB}}(e^{j\frac{2\pi k}{N}})$, respectively. The entry of matrix $\mathbf{V}\in\mathbb{C}^{K\times N}$ in the $k^{\prime}$-th row and $n$-th column is $e^{-j\frac{2\pi nk}{N}}$, where $\frac{2\pi k}{N}$ is the $k^{\prime}$-th discrete frequency in $(\rho\pi+\Omega_{\mathrm{SL}},2\pi-\rho\pi-\Omega_{\mathrm{SL}})$. Let $\underline{\mathbf{z}_{w,\mathcal{S}_{i}}}\in\mathbb{R}^{|\mathcal{S}_{i}|\times 1}$ and $\underline{\boldsymbol{\epsilon}_{w,\mathcal{S}_{i}}}\in\mathbb{R}^{|\mathcal{S}_{i}|\times 1}$ contain the elements of $\underline{\mathbf{z}_{w}}$ and $\underline{\boldsymbol{\epsilon}_{w}}$ whose indices $n\in\mathcal{S}_{i}$. Then, a pre-estimate of $\underline{\mathbf{z}_{w,\mathcal{S}_{i}}}$, denoted as $\underline{\tilde{\mathbf{z}}_{w,\mathcal{S}_{i}}}$, can be expressed as
@@
Note that the $n$-th element of $\underline{\tilde{z}^{(i)}_{w}}[n]$ for $n\in\left\{0,\cdots,\frac{\alpha N}{2}\right\}$ has a non-unity scaling of the desired modulo residue signal $\underline{z^{(i)}}[n]$ due to the tapered cosine window $w[n]$. To correct this scaling, we exploit the symmetry of the window function at the edges:
@@
is applied. The modulo residue signal estimate $\hat{z}^{(i)}[n]$ is then subtracted from $f_{\lambda^{\prime},\mathrm{q}}^{(i)}[n]$ to unfold the $i$-th segment for $n\in\{0,\cdots,N(1-\frac{\alpha N}{2})\}$. Finally, a digital lowpass filter with passband region $\left(-\frac{\Omega_{m}}{2}-\Omega_{\mathrm{SL}},+\frac{\Omega_{m}}{2}+\Omega_{\mathrm{SL}}\right)$ is applied to the result, where $\Omega_{m}=\omega_{m}T_{\mathrm{s}}$. The unfolded signal can be expressed as
@@
for $n\in\{0,\cdots,N(1-\frac{\alpha N}{2})\}$. The subscript $\mathrm{LPF}$ is used to indicate that the modulo residue error signal and quantization noise signal were filtered. The recovery procedure continues until all segments of $f_{\lambda^{\prime},\mathrm{q}}[n]$ have been processed.

### v2

@@ Abstract
The modulo analog-to-digital converter (ADC) is a promising solution to resolve the limited dynamic range (DR) issue of conventional ADCs and achieve an enhanced digital resolution given a fixed quantization bit budget. However, a modulo ADC requires an unfolding scheme to correct the nonlinear distortion introduced by the modulo operation. This paper presents a sliding discrete Fourier Transform (DFT)-based method for fast signal reconstruction given the modulo ADC output sequence and a 1-bit folding information sequence. In contrast to existing DFT-based signal recovery techniques for modulo ADCs, our proposed sliding DFT method reduces the required observation time and minimizes the spectral leakage effects via proper choice of window function parameters. A mean squared error (MSE) performance guarantee is established for the proposed signal recovery algorithm. More precisely, we derive sufficient conditions for the oversampling factor ($\mathrm{OF}$) and the number of quantization bits ($b$) to obtain a specific MSE performance. Our numerical results demonstrate that modulo ADCs equipped with our proposed recovery method can outperform conventional ADCs without modulo for $\mathrm{OF}\geq 4$ and $b\geq 4$. The impact of spectral leakage on the MSE performance of the proposed sliding DFT recovery method is also quantified.
@@
Modulo ADC, Sampling, Quantization, Sliding DFT, Reconstruction.
@@ I Introduction
A promising solution to address the DR bottleneck of ADCs is the modulo sampling framework [6, 7]. The idea in modulo sampling is to apply a nonlinear folding operation called modulo to the input signal before the ADC. This modulo preprocessing stage ensures that the signal range is confined within the DR of the ADC. Moreover, for a given amplitude quantization bit budget, modulo ADCs can achieve better digital resolution compared to conventional ADCs as demonstrated in prior works [8, 9].
@@
Since modulo operation intentionally introduces nonlinear distortion to the signal, several studies on modulo sampling have focused on the development of robust recovery techniques to correct the distortion induced by the modulo operation. Reconstruction techniques for modulo sampling include higher-order differences-based approaches [7], discrete Fourier transform (DFT)-based techniques [10, 11, 12, 13, 14, 15], prediction-based filtering [16, 17], iterative methods [18], sparse signal modulo recovery [19, 20, 21, 22], multi-channel modulo structures [23, 24, 25, 26], and thresholding-based schemes[27, 28, 29, 30, 31]. There are also research efforts geared toward the practical hardware implementation of modulo ADCs [10, 32, 33, 34, 35, 36] and their potential applications in radars, communication systems, imaging, and electroencephalogram (EEG) recovery [37, 38, 39, 40, 12, 41, 42, 43, 25, 44, 45].
@@
Another important research direction in modulo ADC is the establishment of theoretical performance guarantees to reconstruct the unfolded signal from modulo samples. In [7], it is shown that the higher-order differences-based reconstruction approach can perfectly reconstruct a finite-energy bandlimited (BL) signal from its modulo samples if the sampling rate used is at least $2\pi e$ ($\approx 17.07$) times the Nyquist rate. This sufficient condition for the sampling rate increases significantly in the presence of bounded noise [7, Theorem 3]. For finite-energy BL signals, perfect reconstruction is possible for any sampling rate approaching the Nyquist rate from above by using a prediction filter with sufficiently long filter length [16]. This approach does not need any side-information such as knowledge of unfolded samples and auto-correlation function. However, the proposed prediction filter-based approach did not take into account the impact of finite quantization. Theoretical guarantees for Fourier-based reconstruction had been established in [10]. However, this approach suffers from spectral leakage and may require long observation windows. To avoid spectral leakage, recovery methods based on time-domain thresholding can also be used. Such methods work for generalized modulo sampling models which consider the modulo hysterisis and the folding transients. However, recovery guarantees of thresholding-based recovery fail under ideal modulo nonlinearities. This is explained in Section III-E. A recent work on modulo ADCs with 1-bit side information [14] established a mean squared error (MSE) performance guarantee when the oversampling factor (OF) is above three and the number of quantization bits is above three. The 1-bit side information identifies the indices in which folding occured. However, the proofs neglected the spectral leakage effects and the contribution of quantization noise at certain time indices.
@@
In this work, we propose a sliding DFT approach to construct the original input samples from the modulo ADC output samples and an additional 1-bit folding information signal. The recovery procedure uses short-time Fourier transform (STFT) to facilitate successive frame-based unfolding of the modulo samples. To significantly reduce the impact of spectral leakage in STFT computation, a window function with smooth transitions at the edges is multiplied to the frame. A MSE performance guarantee for the proposed recovery method is also established and is compared with the simulated MSE results.
@@
  • We develop an unfolding algorithm for modulo ADC output with 1-bit folding information signal based on sliding-window DFT. In contrast to existing Fourier-based reconstruction techniques, the proposed algorithm does not need a long observation window to process the modulo ADC output. This significantly reduces the computational complexity of the recovery algorithm. Furthermore, spectral leakage due to the short frame observations can be reduced by choosing an appropriate window function. The computational complexity of the proposed recovery algorithm is derived in Section III-G.
  • We establish sufficient conditions for the oversampling factor (OF) and quantizer resolution ($b$) of the modulo ADC so that the proposed recovery algorithm can correctly unfold the modulo ADC output. Under these conditions, only the quantization noise contributes to the MSE. The MSE performance guarantee of the sliding DFT-based recovery is derived in Theorem 1.
@@
Notation: The following notations are used throughout the paper. The sets of real numbers, integers, and natural numbers are denoted as $\mathbb{R}$, $\mathbb{Z}$, and $\mathbb{N}$, respectively. When referring to discrete-time signals, the notation $z[n]$ is used to signify $z(nT_{\mathrm{s}})$, assuming that the sampling period $T_{\mathrm{s}}$ is evident from the context. The first-order difference of a discrete-time signal $x[n]$ is denoted as $\underline{x}[n]=x[n]-x[n-1]$, where $x[-1]=0$ unless stated otherwise. Vectors and matrices are written in bold format (e.g., $\mathbf{z}$, $\mathbf{A}$) while sets are written in calligraphic format (e.g., $\mathcal{S}$). The cardinality of a set $\mathcal{S}$ is written as $|\mathcal{S}|$. If $\mathbf{x}$ is a vectorized form of some discrete-time signal $x[n]$, then $\underline{\mathbf{x}}$ is the vectorized form of $\underline{x}[n]$. The $\ell_{p}$-norm of a vector $\mathbf{x}$ is written as $\|\mathbf{x}\|_{p}$. The set of values of the discrete-time signal $x[n]$ for $n\in\mathcal{S}$ is denoted as $x_{\mathcal{S}}[n]$. The Moore-Penrose inverse of a matrix $\mathbf{A}\in\mathbb{R}^{m\times n}$ is denoted by $\mathbf{A}^{\dagger}$. We use $\omega$ and $\Omega$ to denote the frequencies for the contininuous-time Fourier Transform (CTFT) and discrete-time Fourier Transform (DTFT) spectrum, respectively. We also use the standard Big-O notation $\mathcal{O}(\cdot)$ to describe asymptotic growth rates in this paper.
@@ II-A Signal Acquisition Process
Figure 1: The schematic diagram of the modulo ADC with 1-bit folding information. The 1-bit folding information $c[n]$ is obtained by sampling $c(t)$.
Figure 2: The folding mechanism and folding information generation of the modulo operator block.
We consider the modulo ADC system shown in Figure 1. The input $f(t)$ is a bandlimited signal with (angular) frequency support $\left[-\frac{\omega_{m}}{2},\;+\frac{\omega_{m}}{2}\right]$ and zero mean. This signal is folded by the modulo operator block to produce the modulo signal $f_{\lambda^{\prime}}(t)$. The operation behind the modulo folding mechanism is shown in Figure 2. This feedback structure for the folding operation is adopted from [34]. Comparator 1 triggers a positive value whenever $f_{\lambda^{\prime}}(t)$ crosses $+\lambda$ from below. Similarly, comparator 2 triggers a negative value whenever $f_{\lambda^{\prime}}(t)$ crosses $-\lambda$ from above. These trigger signals are fed to a discrete voltage generator (DVG). The voltage level of the DVG is increased (resp. decreased) by $2\lambda^{\prime}$ whenever it receives a positive (resp. negative) trigger signal. A detailed circuit-level implementation of the DVG is presented in [34]. Mathematically, the folded signal can be expressed as
@@
where $\lambda^{\prime}\in(0,\|f(t)\|_{\infty})$ is the modulo threshold. The $\ell_{\infty}$-norm of a waveform is its maximum amplitude. Throughout this paper, we assume ideal modulo nonlinearities. That is, there are no modulo hysterisis and folding transients like in the generalized modulo sampling model presented in [28].
To digitize the modulo signal, $f_{\lambda^{\prime}}(t)$ is first sampled at every $T_{s}$ seconds. The angular sampling rate is $\omega_{s}=\frac{2\pi}{T_{s}}=\mathrm{OF}\times\omega_{m}$, where $\mathrm{OF}\geq 1$ is the oversampling factor. We also define $\rho=\frac{1}{\mathrm{OF}}$. Prior to quantization, a (non-subtractive) dither sequence $d[n]$ is added to $f_{\lambda^{\prime}}[n]$. Samples of the dither sequence are drawn i.i.d. from a triangle distribution with amplitude support $\big{(}-\frac{2\lambda}{2^{b}},+\frac{2\lambda}{2^{b}}\big{]}$. The rationale for using triangle dither in the modulo ADC problem setup is explained in Section III-B. The resulting signal is then forwarded to the $b$-bit uniform scalar quantizer $\mathcal{Q}_{b}(\cdot)$ with range $[-\lambda,+\lambda]$. Consequently, each quantization bin has width $\frac{2\lambda}{2^{b}}$. To prevent the quantizer from being overloaded, the quantizer DR is set to $\lambda=\frac{2^{b}\lambda^{\prime}}{2^{b}-2}$. The quantizer output can be written as
@@
In addition to $f_{\lambda^{\prime}}(t)$, the modulo operator also generates a signal $c(t)$ according to Figure 2. This signal is a finite rate of innovation (FRI) signal and can be expressed as
@@
where $\tau_{k}$ is the location of the level crossings, $g_{1}(t)$ is the pulse shape of the trigger signals, $g_{2}(t)$ is the sampling kernel, and $g(t)=(g_{2}\ast g_{1})(t)$. To simplify the analysis, $g(t)$ is assumed to be a rectangular pulse with width $T_{s}$ and unity amplitude. The signal $c(t)$ is sampled at rate $\omega_{s}$ and then fed to a 1-bit comparator to produce $c[n]$. In essence, $c[n]$ provides information if the signal crossed either $+\lambda^{\prime}$ or $-\lambda^{\prime}$ in the interval $(nT_{\mathrm{s}},(n+1)T_{\mathrm{s}}]$. We note that even though our modulo ADC generates two channel streams, our setup is different from the multichannel modulo sampling structures presented in [23, 24, 25]. The multiple channels in [23, 24, 25] use different modulo thresholds to process the same signal, thereby generating two multi-bit modulo ADC output sample streams. In contrast, our second stream only indicates the position of the folding instances in the first stream. The information carried by this second stream can be represented by a single bit per sample.
@@ II-B Is the 1-bit Folding Information Practical?
One might wonder if generating 1-bit folding information $c[n]$ is practical when implementing a modulo ADC. Early works on modulo ADC hardware have considered recovery methods that utilize the reset count in addition to the self-reset ADC output samples [46, 47]. However, implementing a self-reset ADC with a multi-bit reset count signal may incur a significant increase in power consumption and circuit complexity if the threshold is small [7]. Furthermore, having a multi-bit reset count signal that has the same resolution as the self-reset ADC output samples doubles the storage requirements of the system. Limiting $c[n]$ to 1-bit solves these problems. As shown in Figures 1 and 2, generating $c[n]$ would only require one additional comparator and one adder from the “no $c[n]$” modulo ADC setup. The additional circuitry has minimal impact on the overall power consumption, and its penalty is fixed regardless of $b$. For instance, consider a flash ADC architecture for amplitude quantization. An 8-bit modulo ADC with no $c[n]$ would have $2582^{8}+2=$ 258 comparators [^1] The 2 additional comparators come from the comparators of the folding mechanism. whereas an 8-bit modulo ADC with 1-bit folding information $c[n]$ would have 259 comparators. The difference in comparator counts is fixed regardless of the number of bits for amplitude quantization. We also note that the folding information generated in Figure 1 is one bit less than the folding information generated by the modulo ADC hardware proposed in [32]. In [32], the 2-bit signal indicates three states: ‘no crossing’, ‘crossed $+\lambda^{\prime}$ threshold’, ‘crossed $-\lambda^{\prime}$ threshold’.
@@ II-C Signal Recovery Scheme
$=$ $w[n]\cdot\left(f_{\lambda^{\prime},\mathrm{q}}^{(i)}[n]-f_{\lambda^{\prime},\mathrm{q}}^{(i)}[n-1]\right),$ (3)
where $f_{\lambda^{\prime},\mathrm{q}}^{(i)}[-1]=f_{\lambda^{\prime},\mathrm{q}}^{(i-1)}[N-1]\;\forall\;i\in\{2,3,\cdots,I\}$ and we assume that $f_{\lambda^{\prime},\mathrm{q}}^{(1)}[-1]=0$, i.e. signal starts at zero. We now describe the computation of the modulo residue samples in $f_{\lambda^{\prime},\mathrm{q}}^{(i)}[n]$. The principle behind the computation of the modulo residue samples is the Fourier domain separation between the bandlimited signal and the folding instances [10]. The length-$N$ Discrete Fourier Transform (DFT) of $\underline{\tilde{f}_{w}^{(i)}}[n]$ is computed as
@@
$\hat{F}_{\mathrm{OOB}}(e^{j\frac{2\pi k}{N}})=$ $\frac{1}{\sqrt{N}}\sum_{n=0}^{N-1}w[n]\bigg{[}\underline{z^{(i)}}[n]+\underline{\epsilon^{(i)}}[n]\bigg{]}e^{-j\frac{2\pi kn}{N}}$
@@
where the spectral leakage width, denoted $\Omega_{\mathrm{SL}}$, accounts for the frequencies where spectral leakage is still significant. Some energy of the $\underline{\tilde{f}_{w}^{(i)}}[n]$ may leak in the band $(\rho\pi,2\pi-\rho\pi)$ due to the inherent spectral leakage in the DFT computation and cannot be neglected. However, spectral leakage can be controlled by properly choosing the window length $N$ and roll-off parameter $\alpha$. The signals $\underline{z^{(i)}}[n]$ and $\underline{\epsilon^{(i)}}[n]$ are the first-order difference of the modulo residue samples and quantization noise, respectively, in $\tilde{f}_{w}^{(i)}[n]$. The set $\mathcal{S}_{i}$ contains the indices for which $c[n]=1$ in the $i$-th segment. The set identifies the indices in the $i$-th segment for which the DVG changes its output voltage. The samples $z^{(i)}_{\mathcal{S}_{i}}[n]$ (resp. $z^{(i)}_{\mathcal{S}_{i}^{c}}[n]$) corresponds to the value of the modulo residue for all $n$ such that $c[n]=1$ (resp. $c[n]=0$) in the $i$-th segment. The subscript $\mathrm{OOB}$ in $\hat{F}_{\mathrm{OOB}}$ is used to indicate that this is the set of DFT coefficients outside the desired signal bandwidth plus non-negligible spectral leakage.
We are primarily interested in estimating $f^{(i)}[n]$. Our approach is to estimate $z^{(i)}[n]$ first and then subtract $z^{(i)}[n]$ from $f_{\lambda^{\prime},\mathrm{q}}^{(i)}[n]$. This procedure essentially unfolds the modulo ADC output. Let $K$ be the number of discrete frequencies $k$ such that $\frac{2\pi k}{N}\in(\rho\pi+\Omega_{\mathrm{SL}},2\pi-\rho\pi-\Omega_{\mathrm{SL}})$. Equation (II-C) can be written in matrix form as
@@
where $\underline{\mathbf{z}_{w}}\in\mathbb{R}^{N\times 1}$, $\underline{\boldsymbol{\epsilon}_{w}}\in\mathbb{R}^{N\times 1}$, and $\hat{\mathbf{F}}_{\mathrm{OOB}}\in\mathbb{C}^{K\times 1}$ are vectorized form of the finite-length signals $w[n]\cdot\underline{z^{(i)}}[n]$, $w[n]\cdot\underline{\epsilon^{(i)}}[n]$, and $\hat{F}_{\mathrm{OOB}}(e^{j\frac{2\pi k}{N}})$, respectively. The entry of matrix $\mathbf{V}\in\mathbb{C}^{K\times N}$ in the $k^{\prime}$-th row and $n$-th column is $e^{-j\frac{2\pi nk^{\prime}}{N}}$, where $\frac{2\pi k^{\prime}}{N}$ is the $k^{\prime}$-th discrete frequency in $(\rho\pi+\Omega_{\mathrm{SL}},2\pi-\rho\pi-\Omega_{\mathrm{SL}})$. Let $\underline{\mathbf{z}_{w,\mathcal{S}_{i}}}\in\mathbb{R}^{|\mathcal{S}_{i}|\times 1}$ and $\underline{\boldsymbol{\epsilon}_{w,\mathcal{S}_{i}}}\in\mathbb{R}^{|\mathcal{S}_{i}|\times 1}$ contain the elements of $\underline{\mathbf{z}_{w}}$ and $\underline{\boldsymbol{\epsilon}_{w}}$ whose indices $n$ are in $\mathcal{S}_{i}$. Then, a pre-estimate of $\underline{\mathbf{z}_{w,\mathcal{S}_{i}}}$, denoted as $\underline{\tilde{\mathbf{z}}_{w,\mathcal{S}_{i}}}$, can be expressed as
@@
Note that the $n$-th element of $\underline{\tilde{z}^{(i)}_{w}}[n]$ for $n\in\left\{0,\cdots,\frac{\alpha N}{2}-1\right\}$ has a non-unity scaling of the desired modulo residue signal $\underline{z^{(i)}}[n]$ due to the tapered cosine window $w[n]$. To correct this scaling, we exploit the symmetry of the window function at the edges:
@@
is applied. The modulo residue signal estimate $\hat{z}^{(i)}[n]$ is then subtracted from $f_{\lambda^{\prime},\mathrm{q}}^{(i)}[n]$ to unfold the $i$-th segment for $n\in\{0,\cdots,N(1-\frac{\alpha N}{2})\}$. The recovery of the unfolded sequence continues until all segments of $f_{\lambda^{\prime},\mathrm{q}}[n]$ have been processed. If $c[n]=0$ for all $n\in\{0,\cdots,N-1\}$ in the $i$-th segment, we can skip the processing of $i$-th segment and proceed to the next one. Finally, a digital lowpass filter with passband region $\left(-\frac{\Omega_{m}}{2}-\Omega_{\mathrm{SL}},+\frac{\Omega_{m}}{2}+\Omega_{\mathrm{SL}}\right)$ is applied to the result, where $\Omega_{m}=\omega_{m}T_{\mathrm{s}}$. The unfolded signal per segment can be expressed as